A boy traveled 10 meters to the east and then he turned 4 meters to the North. What is his resultant displacement.

1
Expert's answer
2022-03-01T11:12:51-0500


The resultant displacement

"d=\sqrt{10^2+4^2}=10.8\:\rm m""\theta=\tan^{-1}\frac{4}{10}=21.8^\circ\;\rm N\: of\; E"